Dungeoness Corporation has excess cash of $1,000 that it would like to distribute to shareholders through a share repurchase. Current earnings are $0.8 per share, and the stock currently sells for $38 per share. There are 280 shares outstanding. Ignore taxes and other imperfections.

If Dungeoness Corp. goes with the share repurchase, what will the price per share be? How many shares will they buy in the repurchase? What are earnings per share (EPS) and the price earnings (P/E) ratio? Enter your answers rounded to 2 DECIMAL PLACES.

Price per share =

Number of shares repurchased =

Earnings per Share =

Price earnings (P/E) ratio =
Answer #1

Price per share = Share Repurchase Price is equal to stock current sells price = $38 per share

Number of shares repurchased = Excess Cash / Share Repurchase Price = 1000/38 = 26.32 Shares

Earnings per Share = Total Earning / No. of Share Outstanding = (0.80 * 280) / (280-26.32) = 224 / 253.68 = $ 0.88

Price earnings (P/E) ratio = Price Per share / Earning Per share = $38/0.88 = 43.18

Please Thumbs Up

Thanks